# SPGD: Steepest Perturbed Gradient Descent Optimization

### v1

@@ Abstract
This paper presents the Steepest Perturbed Gradient Descent (SPGD), a novel algorithm that innovatively combines the principles of the gradient descent method with periodic uniform perturbation sampling to effectively circumvent these impediments and lead to better solutions whenever possible. SPGD is distinctively designed to generate a set of candidate solutions and select the one exhibiting the steepest loss difference relative to the current solution. It enhances the traditional gradient descent approach by integrating a strategic exploration mechanism that significantly increases the likelihood of escaping sub-optimal local minima and navigating complex optimization landscapes effectively. Our approach not only retains the directed efficiency of gradient descent but also leverages the exploratory benefits of stochastic perturbations, thus enabling a more comprehensive search for global optima across diverse problem spaces. We demonstrate the efficacy of SPGD in solving the 3D component packing problem, an NP-hard challenge. Preliminary results show a substantial improvement over four established methods, particularly on response surfaces with complex topographies and in multidimensional non-convex continuous optimization problems. Comparative analyses with established 2D benchmark functions highlight SPGD’s superior performance, showcasing its ability to navigate complex optimization landscapes. These results emphasize SPGD’s potential as a versatile tool for a wide range of optimization problems.
@@ 2 Related Work
Simulated annealing (SA) [4] and genetic algorithm [5] are heuristic optimization algorithms that use randomness and selection mechanisms inspired by natural processes to explore the solution space and select the best candidates for further iteration. These algorithms can be used for different types of optimization problems such as continuous, discrete, non-convex, and multi-objective problems [6, 7]. These methods may require significant computational resources and careful tuning of parameters (e.g., temperature in simulated annealing or mutation rate in evolutionary algorithms) to balance exploration and exploitation effectively.
@@
Nesterov’s Accelerated Gradient (NAG) method enhances traditional gradient descent by incorporating a forward-looking step. This tweak allows the optimizer to anticipate future gradients, reducing oscillations and speeding up convergence, particularly in convex settings. NAG is highly effective in training deep neural networks due to its efficiency in navigating high-dimensional data spaces. However, its performance can vary in non-convex environments with complex landscapes [8, 9].
@@ 3 SPGD
To demonstrate the execution flow of the SPGD algorithm and to provide intuitive insights into its operation, Figure 1 offers a detailed visualization of a simple 1D parabolic landscape. This example illustrates the iterative process where traditional gradient descent, represented by the green line, begins the optimization journey. Upon activation of the perturbation process after iterations, SPGD generates candidates (depicted as blue pluses) around the current position. These candidates explore the vicinity of the solution space, identifying potential points with lower function values. The most promising candidate, offering the lowest function value, is then compared to the current position. As depicted, this candidate yields a better function value and is thus chosen, causing the algorithm to make a significant ”jump” to this new point. The process then resumes with gradient descent steps leading to faster convergence to the optimal solution. This example effectively showcases how SPGD combines exploration and exploitation to enhance convergence speed in scenarios where traditional gradient descent may progress slowly.
@@ 4 Numerical Results
We present here a thorough evaluation of the proposed Steepest Perturbed Gradient Descent (SPGD) algorithm, comparing its performance against several established optimization methods. The comparison includes traditional gradient descent (GD), Perturbed Gradient Descent (PGD), MATLAB function which is a versatile solver for constrained optimization problems [16], and Simulated Annealing (SA) [17].
Our initial analysis is conducted through the lens of four challenging 2D benchmark functions, selected for their known difficulties and relevance in assessing optimization algorithms’ efficacy. These test functions are recognized benchmarks within the optimization community, providing a diverse set of landscapes to evaluate each algorithm’s ability to navigate complex, non-convex, and potentially deceptive optimization spaces [18]. For each test function, we apply , Simulated Annealing, traditional gradient descent, PGD, and SPGD, meticulously recording and analyzing the results. In both the SPGD and PGD algorithms, the amplitude of the perturbation is set to be the same, ensuring comparability, is set to , and is set to . Furthermore, the gradient function is explicitly passed to the function in each scenario to guide the optimization process.
Key performance indicators include the accuracy of the solution, measured by the proximity to the known global optimum[19]; the computational efficiency, quantified by the number of function evaluations and CPU execution time. The SPGD algorithm’s source code, along with comparative analyses against the discussed methods using additional 2D challenging test functions, is publicly accessible on GitHub 333Source code and comparisons available at: https://github.com/Amir-M-Vahedi/SPGD-Benchmark-Functions.
@@
It has a global minimum point located at , with an optimal function value of . The initial condition is chosen randomly to be , and the is set to 2.5. Figure 2 and 3 illustrate the 3D view and top view of optimization trajectory across the Peaks function surface. The convergence history comparison is shown in Fig 4. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 1. Based on the results depicted in Figures 2, 3, and 4, and performance metrics in Table 1, it is evident that the GD and PGD algorithms become trapped in local minima. In contrast, the , SA, and SPGD algorithms successfully converge to the global optimum. Among these three, SPGD demonstrates the lowest computational cost. Notably, despite the method having fewer function evaluations, its CPU time is approximately 20 times greater than that of the SPGD algorithm.
@@
The initial condition is chosen randomly to be , and the is set to 2.5. Figure 5 and 6 illustrate the 3D view and top view of optimization trajectory across the Ackley function surface. The convergence history comparison is shown in Fig 7. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 2. Taking into account the data presented in the mentioned figures and table, analysis of the Ackley test function reveals that the GD, PGD, and methods became ensnared in local minima. In contrast, only the SA and SPGD algorithms successfully navigated to the global solution. Among these, SPGD not only achieved convergence with greater precision—approaching closer to the global optimum—but also demonstrated a computational speed, with CPU execution time being six times faster than its counterpart.
@@
The initial condition is chosen randomly to be , and the is set to 5. Figure 8 and 9 illustrate the 3D view and top view of optimization trajectory across the Easom function surface. The convergence history comparison is shown in Fig 10. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 3. Reflecting on the performance metrics for the Easom test function, it is evident that only the SPGD algorithm successfully pinpointed the global optimum. As anticipated, GD was hindered in its progression by the minimal gradient values inherent to the function’s extensive flat regions. Similarly, both GD and the method failed to escape these flat expanses, effectively becoming ensnared within them. Among the competing methods, only SA managed to navigate towards a more favorable outcome, yet it fell short of achieving convergence to the global optimum, underscoring the distinctive effectiveness of SPGD in this scenario.
@@
The initial condition is chosen randomly to be , and the is set to 2.5. Figure 11 and 12 illustrate the 3D view and top view of optimization trajectory across the Levy Function No. 13 surface. The convergence history comparison is shown in Fig 13. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 4. Based on the performance analysis for this test function, the GD, PGD, and methods were unable to find the global optimum, getting stuck in local minima instead. Notably, settled in a particularly poor local minimum. In contrast, both Simulated Annealing (SA) and SPGD successfully navigated to the global optimum. However, SPGD distinguished itself by achieving a more accurate solution, requiring fewer function evaluations, and demonstrating a 25 times faster CPU execution time compared to SA.
@@
| GD | 2001 | 6.2915 | *4.58 |
| PGD | 2001 | 6.2915 | *3.85 |
| 20 | 30.5009 | *88.90 | |
| SA | 2018 | 6.78-7 | 175.38 |
| SPGD | 1760 | 2.45-13 | 7.56 |
@@
Our 3D packing scenarios presented here are specialized instances of those tackled by SPI2-F – a novel and more general packing and layout optimization presented in [34], that performs both packing and layout optimization of complex interconnected systems in a multi-physics environment. The specialized scenarios presented below have been chosen to include cases that have known global optima. We note that an efficient packing method based on Fast Fourier Transforms444Fast Fourier Transforms have been previously shown to offer an elegant and efficient approach to compute collisions and penetrations as well as shape complementarity [35, 36, 37]. was introduced recently in [38], which restricts the orientation of the objects to an axis-alignment and hence allows rotations in 90-degree increments. By contrast, our method allows arbitrary rotations and alignments in space.
@@ 4.4 Analysis of Scenario 2: Eight Identical Cubes
In Scenario 2, the initial configuration of eight identical cubes is depicted in Figure 18. This scenario was designed to assess each algorithm’s ability to scale and manage increased numbers of objects while maintaining an efficient packing configuration. The outcomes of the final configurations found by the GD and SPGD algorithms are illustrated in Figures 19(a) and 19(b), respectively.
@@ 4.5 Analysis of Scenario 3: Eight Cubes of Different Sizes
The SPGD algorithm’s performance in this scenario was notably superior, as it converged to a more compact solution significantly faster than the traditional GD method. The results of the final configurations found by the GD and SPGD algorithms are shown in Figures 22(a) and 22(b), respectively.
@@ 4.6 Analysis of Scenario 4: Eight Objects of Different Shapes
In this demanding scenario, the SPGD algorithm demonstrated its robust capability by converging to a significantly more compact solution compared to the traditional GD method. Although the time taken by SPGD to find the optimal solution was comparable to that of GD, the overall optimization process required more time due to the termination condition set for no improvement in the loss value over 2000 iterations. The final configurations achieved by the GD and SPGD algorithms are shown in Figure 25(a) and 25(b), reflecting the SPGD algorithm’s effectiveness in handling complex and varied object forms.

### v2

@@ Abstract
This paper presents the Steepest Perturbed Gradient Descent (SPGD), a novel algorithm that innovatively combines the principles of the gradient descent method with periodic uniform perturbation sampling to effectively circumvent these impediments and lead to better solutions whenever possible. SPGD is distinctively designed to generate a set of candidate solutions and select the one exhibiting the steepest loss difference relative to the current solution. It enhances the traditional gradient descent approach by integrating a strategic exploration mechanism that significantly increases the likelihood of escaping sub-optimal local minima and navigating complex optimization landscapes effectively. Our approach not only retains the directed efficiency of gradient descent but also leverages the exploratory benefits of stochastic perturbations, thus enabling a more comprehensive search for global optima across diverse problem spaces. We demonstrate the efficacy of SPGD in solving the 3D component packing problem, an NP-hard challenge. Preliminary results show a substantial improvement over six established methods, particularly on response surfaces with complex topographies and in multidimensional non-convex continuous optimization problems. Comparative analyses with established 2D benchmark functions over 30 randomized initial points highlight SPGD’s robustness and reliability in non-convex optimization. These results emphasize SPGD’s potential as a versatile tool for a wide range of optimization problems.
@@ 2 Related Work
Simulated annealing (SA) [4] and genetic algorithm [5] are heuristic sampling-based optimization algorithms that use randomness and selection mechanisms inspired by natural processes to explore the solution space and select the best candidates for further iteration. These algorithms can be used for different types of optimization problems, such as continuous, discrete, non-convex, and multi-objective problems [6, 7]. These methods may require significant computational resources and careful tuning of parameters (e.g., temperature in simulated annealing or mutation rate in evolutionary algorithms) to balance exploration and exploitation effectively.
Bayesian optimization (BO) is one of the sampling-based global optimization methods that has gained popularity, particularly in machine learning, for solving expensive black-box optimization problems. BO methods approximate the objective function using a surrogate probabilistic model, typically a Gaussian process (GP), which models the underlying function based on observed sample points [8, 9]. These methods balance exploration and exploitation by combining prior beliefs with posterior updates after each observation. The acquisition function, derived from this surrogate model, guides the search by quantifying the expected improvement or uncertainty in unexplored regions. Although BO is highly sample-efficient and effective in finding global optima with a limited number of function evaluations, especially in low-dimensional problems, it can be computationally expensive due to the cost of updating and optimizing the acquisition function at each iteration. Moreover, the performance of BO degrades in high-dimensional or highly non-smooth optimization landscapes [10]. A broader discussion of such sampling-based approaches can be found in [11], a recent survey on non-smooth optimization methods, including gradient sampling and probabilistic techniques.
@@
Nesterov’s Accelerated Gradient (NAG) method enhances traditional gradient descent by incorporating a forward-looking step. This tweak allows the optimizer to anticipate future gradients, reducing oscillations and speeding up convergence, particularly in convex settings. NAG is highly effective in training deep neural networks due to its efficiency in navigating high-dimensional data spaces. However, its performance can vary in non-convex environments with complex landscapes [12, 13]. For a comprehensive overview of gradient descent and its variants, we refer readers to [2], which synthesizes developments across machine learning and optimization literature.
@@ 3 Methodology: SPGD
To demonstrate the execution flow of the SPGD algorithm, we provide a visual comparison with traditional gradient descent (GD) in Figure 1, applied to a non-convex function defined by . This landscape features both a local minimum and a global minimum, offering an ideal setting to showcase the strengths of SPGD in escaping poor regions of convergence. In figure 1, GD is seen following the steepest descent path, ultimately settling at the local minimum without the ability to recover. This behavior is the characteristic of gradient-based methods in non-convex landscapes, where they are prone to getting trapped due to the absence of global information or exploration strategies.
The SPGD algorithm, on the other hand, begins similarly by following the gradient descent path, represented by the green line, for a fixed number of iterations. After iterations, the perturbation phase is triggered. At this point, SPGD generates candidates (depicted as black ’x’) around the current position. These candidates explore the vicinity of the solution space, identifying potential points with lower function values. The most promising candidate, offering the lowest function value (shown as blue ’+’), is then compared to the current position. As depicted, this candidate yields a better function value and is thus chosen, causing the algorithm to make a significant ”jump” to this new point. In the figure 1, this is reflected in the trajectory of SPGD deviating from the region of slow progress of local minimum and moving toward the global one. The process then resumes with gradient descent steps leading to faster convergence to the optimal solution. As a result, SPGD not only avoids becoming stuck in local minima but also achieves convergence with fewer function evaluations compared to methods that either rely solely on gradients or purely stochastic exploration. This example highlights how SPGD’s structure, consisting of gradient-based updates interleaved with perturbations, contributes to both its robustness and efficiency in solving non-convex optimization problems.
@@ 4 Numerical Results
We present here a thorough evaluation of the proposed Steepest Perturbed Gradient Descent (SPGD) algorithm, comparing its performance against several established optimization methods. The comparison includes traditional gradient descent (GD), Perturbed Gradient Descent (PGD), MATLAB function, which is a versatile solver for constrained optimization problems [20], and the function, which is tailored to unconstrained optimization problems [21], along with Simulated Annealing (SA) [22], and Bayesian Optimization (BO) [23].
Our initial analysis is conducted through the lens of four challenging 2D benchmark functions, selected for their known difficulties and relevance in assessing optimization algorithms’ efficacy. These test functions are recognized benchmarks within the optimization community, providing a diverse set of landscapes to evaluate each algorithm’s ability to navigate complex, non-convex, and potentially deceptive optimization spaces [24]. For each test function, we apply , Simulated Annealing, traditional gradient descent, PGD, and SPGD, meticulously recording and analyzing the results. The SPGD and PGD algorithms were each fine-tuned independently to ensure optimal performance while maintaining a fair basis for comparison. Due to the high computational cost of Bayesian Optimization (BO), the maximum number of function evaluations for BO is capped at 100 to ensure reasonable execution time across benchmark functions. For the function, we use MATLAB’s default interior-point algorithm, while the function is configured to use the trust-region method, which is well-suited for smooth, unconstrained problems. In both cases, the gradient of the objective function is explicitly provided to guide the optimization process more efficiently.
Key performance indicators include the accuracy of the solution, measured by the proximity to the known global optimum[25]; the computational efficiency, quantified by the number of function evaluations and CPU execution time. For each test function, both a 3D and top-view surface plot of optimization trajectory visualization are provided to aid in understanding each algorithm’s optimization landscape and behavior. These visualizations illustrate how optimization paths evolve over complex response surfaces and help highlight differences in convergence dynamics. Simulated Annealing (SA) and Bayesian Optimization (BO) are excluded from these visualizations, as their probabilistic sampling strategies tend to densely populate the landscape, obscuring the trajectories of other algorithms and reducing the overall interpretability of the plots. The source code for the SPGD algorithm, along with comparative analyses against methods discussed in this paper using additional 2D challenging test functions, are publicly accessible on GitHub333Source code and comparisons available at: https://github.com/Amir-M-Vahedi/SPGD-Benchmark-Functions.
@@
It has a global minimum point located at , with an optimal function value of . The initial condition is chosen randomly to be , and the is set to 2.5. Figure 2(a) and 2(b) illustrate the 3D view and top view of the optimization trajectory across the Peaks function surface. The total number of function evaluations, the converged optimal value, and CPU execution time for different methods are given in Table 1. Based on the results depicted in Figures 2, and performance metrics in Table 1, it is evident that the GD, PGD, and algorithms become trapped in local minima. In contrast, the , SA, BO, and SPGD algorithms successfully converge to the global optimum. Among these three, SPGD demonstrates the lowest computational cost. Notably, despite the and BO method having fewer function evaluations, their CPU times are more than 25 and 2000 times greater than that of the SPGD algorithm.
@@
The initial condition is chosen randomly to be , and the is set to 2.5. Figure 3(a) and 3(b) illustrate the 3D view and top view of optimization trajectory across the Ackley function surface. The performance comparisons are given in Table 2. Taking into account the data presented in the mentioned figures and table, analysis of the Ackley test function reveals that the GD, PGD, , and methods became ensnared in local minima. In contrast, only the SA, BO, and SPGD algorithms successfully navigated to the global solution. Among these, SPGD not only achieved convergence with greater precision, approaching closer to the global optimum, but also demonstrated a computational speed, with the CPU execution time being about 13 (SA) and 359 (BO) times faster than its counterparts.
@@
The initial condition is chosen randomly to be , and the is set to 5. Figure 4(a) and 4(b) illustrate the 3D view and top view of the optimization trajectory across the Easom function surface. The performance comparisons are given in Table 3. Reflecting on the performance metrics for the Easom test function, it is evident that only the SPGD algorithm successfully pinpointed the global optimum. As anticipated, GD was hindered in its progression by the minimal gradient values inherent to the function’s extensive flat regions. Similarly, both GD and the method failed to escape these flat expanses, effectively becoming ensnared within them. Among the competing methods, only SA and BO managed to navigate towards a more favorable outcome, yet they fell short of achieving convergence to the global optimum, underscoring the distinctive effectiveness of SPGD in this scenario.
@@
The initial condition is chosen randomly to be , and the is set to 2.5. Figure 5(a) and 5(b) illustrate the 3D view and top view of the optimization trajectory across the Levy Function No. 13 surface. The performance comparisons are given in Table 4. Based on the performance analysis for this test function, the GD, PGD, , and methods were unable to find the global optimum, getting stuck in local minima instead. Notably, settled in a particularly poor local minimum. In contrast, SA, BO, and SPGD successfully navigated to the global optimum. However, SPGD distinguished itself by achieving a more accurate solution, requiring fewer function evaluations than SA, and demonstrating faster CPU execution time compared to SA and BO.
@@
| GD | 2001 | 6.2915 | *3.58 |
| PGD | 2001 | 6.2915 | *2.58 |
| 9 | 14.3717 | *20.59 | |
| 20 | 30.5009 | *53.32 | |
| SA | 2018 | 6.78-7 | 89.61 |
| BO | 100 | 0.0086 | 5241.7 |
| SPGD | 1760 | 2.45-13 | 5.02 |
Robustness Evaluation
To evaluate the robustness of the SPGD algorithm, each test function was subjected to 30 independent trials using randomly sampled starting points. All optimization algorithms were provided with the same lower and upper bounds defining the feasible search space. Each algorithm was fine-tuned independently to ensure their optimal individual performance, and consistent parameter settings were applied across all trials to maintain fairness. The following performance criteria are reported in Tables 5–8: the number of successful convergences (ConvergedRuns), the average percentage improvement in objective value and CPU execution time relative to SPGD (Fval Improvement%, Time Improvement%), and the average closeness to the global optimum (Closer%), over 30 randomized runs, which reflects how much nearer each method’s average result is to the optimum compared to SPGD. In cases where SPGD has already achieved the global optimum or the baseline method matched it, this value is marked as “N/A”.
For the Peaks function, SPGD successfully converged to the global optimum in all 30 trials. Only Bayesian Optimization (BO) matched this convergence count, with Simulated Annealing (SA) achieving 29 out of 30. However, SPGD accomplished this with significantly lower computational cost, as reflected in the Time Improvement% column of Table 5, demonstrating superior efficiency.
In the case of the Ackley function, SPGD was the only algorithm to consistently converge to the global solution across all runs. Although GD and PGD had lower average execution times, they only succeeded in 3 and 5 out of 30 runs respectively, making them less competitive. SPGD outperformed all remaining methods in terms of average speed and reliability over 30 trials.
For the Easom function, none of the baseline algorithms found the global optimum in any run. SPGD was the only method to successfully reach the global solution in all 30 trials, highlighting its robustness in highly deceptive landscapes.
In the Levi function N.13, SPGD again demonstrated the highest reliability with 30 successful runs out of 30. BO and SA achieved 24 successful runs each, while other algorithms failed to find the global optimum in any trial. SPGD also demonstrated strong efficiency, with an average Time Improvement% of 99.91 over BO and 90.02 over SA.
| Algorithm | ConvergedRuns | Fval Improvement % | Time Improvement % | Closer % |
| GD | 9 | 135.63 | -5.43 | 100.00 |
| PGD | 9 | 127.32 | -10.01 | 100.00 |
| BayesOpt | 30 | 0.00 | 99.99 | N/A |
| SA | 29 | 0.04 | 98.90 | 98.71 |
| Fminunc | 7 | 206.33 | 81.57 | 100.00 |
| Fmincon | 8 | 137.05 | 93.10 | 100.00 |
| SPGD | 30 |
| Algorithm | ConvergedRuns | Fval Improvement % | Time Improvement % | Closer % |
| GD | 3 | 99.97 | -1240.80 | 99.97 |
| PGD | 5 | 99.97 | -1079.38 | 99.97 |
| BayesOpt | 17 | 83.04 | 99.96 | 83.04 |
| SA | 27 | 99.15 | 85.82 | 99.15 |
| Fminunc | 5 | 99.97 | 50.29 | 99.97 |
| Fmincon | 8 | 99.96 | 83.92 | 99.96 |
| SPGD | 30 |
| Algorithm | ConvergedRuns | Fval Improvement % | Time Improvement % | Closer % |
| GD | 0 | N/A | -79148.81 | 100.00 |
| PGD | 0 | 444556889.89 | -1822.73 | 100.00 |
| BayesOpt | 0 | 4137.02 | 99.97 | 100.00 |
| SA | 0 | 8841.39 | 78.76 | 100.00 |
| Fminunc | 0 | N/A | -724.67 | 100.00 |
| Fmincon | 0 | N/A | -277.97 | 100.00 |
| SPGD | 30 |
| Algorithm | ConvergedRuns | Fval Improvement % | Time Improvement % | Closer % |
| GD | 0 | 100.00 | -1761.37 | 100.00 |
| PGD | 0 | 100.00 | -1462.19 | 100.00 |
| BayesOpt | 24 | 100.00 | 99.91 | 100.00 |
| SA | 24 | 100.00 | 90.02 | 100.00 |
| Fminunc | 0 | 100.00 | -140.21 | 100.00 |
| Fmincon | 0 | 100.00 | 34.57 | 100.00 |
| SPGD | 30 |
@@
Our 3D packing scenarios presented here are specialized instances of those tackled by SPI2-F – a novel and more general packing and layout optimization presented in [40], that performs both packing and layout optimization of complex interconnected systems in a multi-physics environment. The specialized scenarios presented below have been chosen to include cases that have known global optima. We note that an efficient packing method based on Fast Fourier Transforms444Fast Fourier Transforms have been previously shown to offer an elegant and efficient approach to compute collisions and penetrations as well as shape complementarity [41, 42, 43, 44]. was introduced recently in [45], which restricts the orientation of the objects to an axis-alignment and hence allows rotations in 90-degree increments. By contrast, our method allows arbitrary rotations and alignments in space.
@@ 4.4 Analysis of Scenario 2: Eight Identical Cubes
In Scenario 2, the initial configuration of eight identical cubes is depicted in Figure 10. This scenario was designed to assess each algorithm’s ability to scale and manage increased numbers of objects while maintaining an efficient packing configuration. The outcomes of the final configurations found by the GD and SPGD algorithms are illustrated in Figures 11(a) and 11(b), respectively.
@@ 4.5 Analysis of Scenario 3: Eight Cubes of Different Sizes
The SPGD algorithm’s performance in this scenario was notably superior, as it converged to a more compact solution significantly faster than the traditional GD method. The results of the final configurations found by the GD and SPGD algorithms are shown in Figures 14(a) and 14(b), respectively.
@@ 4.6 Analysis of Scenario 4: Eight Objects of Different Shapes
In this demanding scenario, the SPGD algorithm demonstrated its robust capability by converging to a significantly more compact solution compared to the traditional GD method. Although the time taken by SPGD to find the optimal solution was comparable to that of GD, the overall optimization process required more time due to the termination condition set for no improvement in the loss value over 2000 iterations. The final configurations achieved by the GD and SPGD algorithms are shown in Figure 17(a) and 17(b), reflecting the SPGD algorithm’s effectiveness in handling complex and varied object forms.